OpenStudy (anonymous):

In ∆ABC, if a = 8.75 centimeters, c = 4.26 centimeters, and m B = 87°, what is the length of b to two decimal places? 8.92 centimeters 9.41 centimeters 9.53 centimeters 13.01 centimeters 15.36 centimeters @lovepurple @StudyGurl14

OpenStudy (anonymous):

yeah do i pick 2 numbers

OpenStudy (anonymous):

im not sure my teacher said he had a hard time with this

OpenStudy (anonymous):

hmm????

OpenStudy (studygurl14):

hold on a sec while i solve

OpenStudy (anonymous):

gotchu

OpenStudy (studygurl14):

Odd...this seems like a difficult problem for you...Anyway...use the law of cosines

OpenStudy (anonymous):

im sorry but im so far lost im in orbit

OpenStudy (anonymous):

i dont get anything

OpenStudy (anonymous):

and i still dont get the answer lol

OpenStudy (anonymous):

i know huh

OpenStudy (anonymous):

13.01 i think

OpenStudy (anonymous):

13.01

OpenStudy (studygurl14):

No @lovepurple approx 9.53

OpenStudy (studygurl14):

\(\large b^2=(8.75)^2+(4.26)^2-2(8.75)(4.26)(cos(87))\) \(\large b=\sqrt{76.5625+18.1476-74.55cos(87)}\) \(\large b\approx 9.529347011...\approx9.53\)

OpenStudy (anonymous):

so 9.53

OpenStudy (anonymous):

k